Quiz Chapter 8 Fernando Orellana submitted Oct 20 at 9:39am O 0.30 O 1.50 O 10.0 O cannot answer without additional information Incorrect 0/2 pts uestion A treatment is administered to a sample selected from a population with a mean of 80 and a standard deviation of σ = 10. After treatment, the effect size is measured by computing Cohens d, and a value of d 0.60 is obtained. Based on this information the mean for the treated sample is M- cess as vas O 86 O cannot be determined without additional information rse FAQS LAN Question 19 2/2 pts The power of a statistical test is the probability of 0 rejecting a true null hypothesis O supporting true null hypothesis rejecting a false null hypothesis me anything

which one is correct ? this was confusing . B is wrong or it is C? chnage the sample size from n=100 to n =25?
Answer #1

Ans:

Cohen's d effect size:

d=\dfrac{M-\mu }{\sigma }

0.6=(M-80)/10

M=80+6

M=86

Correct option is 86